OpenStudy (rock_mit182):

A chain store manager has been told by the main office that daily profit, P, is related to the number of clerks working that day, x, according to the function P = −25x^2 + 300x. What number of clerks will maximize the profit, and what is the maximum possible profit?

OpenStudy (rock_mit182):

please help me out i don't know how to find, without derivatives

OpenStudy (rock_mit182):

@ganeshie8

ganeshie8 (ganeshie8):

any idea on how the graph of given function looks like ?

ganeshie8 (ganeshie8):



OpenStudy (rock_mit182):

yeah sure parabola downwards

OpenStudy (rock_mit182):

ohh i see the vertex

ganeshie8 (ganeshie8):

yes, just find the vertex

OpenStudy (rock_mit182):

but how ?

OpenStudy (rock_mit182):

can you show me?, please

ganeshie8 (ganeshie8):

\(\large f(x) = ax^2 + bx+c\) x coordinate of vertex = \(\large \dfrac{-b}{2a}\)

OpenStudy (rock_mit182):

ohh ok ...o Demonstrate your solution for both the number of clerks that will maximize the profit as well as the total maximum profit possible, making sure to include all mathematical work and an explanation for each step.

OpenStudy (rock_mit182):

so in order to solve the last part, i just have to show all the steps finding the vertex

OpenStudy (rock_mit182):

?

OpenStudy (rock_mit182):